12 – 18

Class Boundaries

__________________________________________

Midpoint

__________________________________________

Width


1
Expert's answer
2021-02-24T06:10:42-0500

Lower Class Boundary: The lower class boundary is the average of the lower limit of that class and the upper limit of previous class.

Upper Class Boundary: The upper class boundary of a class is the average of the upper limit of that class and lower limit of the next class.

Mid Point: The mid-point is the average of the class-boundaries of that class.

Width: The difference between upper and lower boundaries of a class is called width of that class.

Class Boundaries:

"11.5- \frac{18+56}{2} = 11.5- 37"

Midpoints:

"\frac{11.5+37}{2}=24.25"

Widths:

"37 -11.5 = 25.5"